OpenStudy (anonymous):

Identify the 16th term of a geometric sequence where a1=4 and a8=-8,748

OpenStudy (anonymous):

a = 4 a_8 = ar^7 = -8748 ==> r^7 = -8748/4 = -2187 r = (-2187)^1/7 = -3 a_16 = ar^15 = 4(-3)^15

OpenStudy (anonymous):

=-57395628